OpenStudy (anonymous):

Given that the straight line y=3x+c is a tangent to the curve y=x^2+9x+k, express k in terms of c.

sam (.sam.):

 intersection y=x^2+9x+k , y=3x+c x^2+9x+k=3x+c rearrange x^2+6x+k-c=0 b^2-4ac=0 a=1 b=6 c=k-c (6)^2-4(1)(k-c)=0 solve

OpenStudy (anonymous):

k=c+18

OpenStudy (anonymous):

Sam : Thanks again (: Khalil : Where did get 18 from? (:

OpenStudy (anonymous):

first u should solve y'(x)=3 let a the solution then y(a)=c

OpenStudy (anonymous):

the line y=3x+c has slope 3 so it must be tangent to y=x^2 + 9x + k at x=-3: 3 = 2x + 9 ---> x=-3 so equating the two equations: 3x + c = x^2 + 9x + k 3(-3) + c = (-3)^2 + 9(-3) + k -9 + c = 9 - 27 + k -9 + c = -18 + k 9 + c = k

OpenStudy (anonymous):

i think we should equat the equations: y(-3)=c